Suppose there are no taxes. Firm ABC has no​ debt, and firm XYZ has debt of $ 6 comma 000 on which it pays interest of 10 % each year. Both companies have identical projects that generate free cash flows of $ 6 comma 100 or $ 6 comma 200 each year. After paying any interest on​ debt, both companies use all remaining free cash flows to pay dividends each year. a. In the table​ below, fill in the debt payments for each firm and the dividend payments the equity holders of each firm will receive given each of the two possible levels of free cash flows. b. Suppose you hold 10 % of the equity of ABC. What is another portfolio you could hold that would provide the same cash​ flows?   c. Suppose you hold 10 % of the equity of XYZ. If you can borrow at 10 % ​, what is an alternative strategy that would provide the same cash​ flows? a. In the table​ below, fill in the debt payments for each firm and the dividend payments the equity holders of each firm will receive given each of the two possible levels of free cash flows.  ​(Round all answers to the nearest​ dollar.)
Answer #1

a) interest = 6000 x 10% = 600

dividends = free cash flows - interest

  XYZ ABC Debt Equity Debt Equity Free cash flows payments Dividends payments Dividends 6100 6100 6001 5500 6200 6200 6001 5600

b)

if you held 10% of equity in ABC then you get 6100 x 10% = 610 as dividend

unlevered equity = levered equity + debt

in order to get same cash flow one should hold 10% of equity and 10% of Debt in XYZ

equity dividend = 5500 x 10% = 550

interest on debt = 600 x 10% = 60 ( total debt = 6000 ,10% would be 6000 x 10% = 600)

total cash flow =  equity dividend +  interest on debt

= 550 + 60

= 610

(same concept for cash flow of 6200 i.e., 620 = 560 + 60 )

c)

levered equity = Borrowing + Unlevered equity

if you have 10% of equity of XYZ you get 5500 x 10% = 550

so one should borrow $600 and buy 10% of equity in ABC

cash flow in ABC = 6100 x 10% = 610

repayment of interest on borrowings = 600 x 10% = 60

so total cash flow = 610 - 60 = $550

(same concept for cash flow of 6200 i.e., 560 = 620 - 60)